# MathCAMPS: Fine-grained Synthesis of Mathematical Problems From Human Curricula

### v1

@@ 3.3 Generating follow-up questions
As human instructors know, follow-up questions are often the best way to probe a student’s understanding. In MathCAMPS, we leverage our symbolic representation of problems to derive follow-up questions. We propose two kinds of questions: counterfactual questions, where we change a constant in the original problem, and incremental questions, where we add a new piece of information. For each CC standard, we mark which (if any) of these two categories of follow-ups are applicable. Symbolically, follow-up questions are represented as a difference to be applied to the original question — when we apply the difference, we obtain a new problem. We then use the same solver as the original problem to obtain the ground-truth answer to the follow-up question. We employ the same few-shot structure to translate this difference into a natural language question, and parse it back into a symbolic structure to test for cycle consistency.
4 Experiments
We now evaluate a suite of 23 LLMs from 8 different vendors on MathCAMPS. We evaluate all models by sampling with temperature 0, using a fixed 1-shot prompt with the first example from GSM8K, mostly to demonstrate the format. For all models (most of them instruction-tuned), a single example was enough for to adhere to the task and the format we specify. The rich structure in MathCAMPS allows us to perform a number of unique analyses on LLMs relating to specific mathematical abilities and their corresponding grade levels for human students. Precisely, we investigate: (1)How do LLMs perform overall on MathCAMPS? How does their performance correlate with GSM8k? (2) Do individual models have relative strengths and weaknesses, or does performance improve uniformly across skills? (3) How well do LLMs respond to follow-up questions? How is their accuracy affected when also considering follow-ups?
4.1 Overall performance
Vendor Model All K 1 2 3 4 5 6 7 8
OpenAI GPT-4o [1] 0.92 0.98 0.98 0.98 0.98 0.92 0.88 0.95 0.89 0.64
Anthropic Claude-3 Opus [2] 0.89 0.97 0.99 0.96 0.98 0.89 0.83 0.96 0.73 0.56
Google Gemini-1.5 Pro [23] 0.89 0.95 0.98 0.97 0.97 0.89 0.83 0.93 0.78 0.54
Google Gemini-1.5 Flash [23] 0.87 0.98 0.98 0.97 0.98 0.80 0.80 0.90 0.84 0.56
OpenAI GPT-3.5 Turbo [1] 0.87 0.96 0.98 0.98 0.97 0.86 0.77 0.90 0.77 0.56
Anthropic Claude-3 Sonnet [2] 0.86 0.96 0.98 0.97 0.98 0.88 0.74 0.94 0.66 0.49
Anthropic Claude-3 Haiku [2] 0.84 0.97 0.98 0.97 0.98 0.87 0.69 0.92 0.59 0.51
Meta Llama 3 70B [25] 0.85 0.96 0.97 0.97 0.97 0.85 0.71 0.87 0.73 0.50
Mistral Mixtral 8x22B [16] 0.84 0.96 0.99 0.98 0.96 0.79 0.69 0.88 0.73 0.61
DeepSeek DeepSeek 67B [5] 0.80 0.95 0.99 0.96 0.93 0.82 0.60 0.84 0.61 0.47
Meta Llama 3 8B [25] 0.77 0.94 0.97 0.96 0.94 0.78 0.55 0.79 0.53 0.43
Mistral Mixtral 8x7B [16] 0.76 0.94 0.96 0.93 0.91 0.75 0.52 0.80 0.53 0.45
EleutherAI Llemma 34B [3] 0.71 0.95 0.96 0.93 0.87 0.61 0.47 0.77 0.46 0.44
Mistral Mistral 7B [15] 0.68 0.89 0.94 0.91 0.84 0.61 0.42 0.66 0.45 0.42
DeepSeek DeepSeek Coder 33B [11] 0.65 0.88 0.93 0.92 0.83 0.54 0.36 0.66 0.44 0.38
Meta CodeLlama 34B [21] 0.64 0.90 0.94 0.92 0.85 0.51 0.38 0.70 0.37 0.30
Microsoft phi-2 [17] 0.63 0.95 0.96 0.89 0.78 0.46 0.38 0.61 0.37 0.41
EleutherAI Llemma 7B [3] 0.62 0.88 0.90 0.85 0.79 0.48 0.41 0.67 0.41 0.36
Google Gemma 7B [24] 0.62 0.83 0.92 0.90 0.82 0.47 0.36 0.65 0.36 0.30
Meta CodeLlama 13B [21] 0.58 0.87 0.92 0.87 0.75 0.41 0.30 0.61 0.32 0.34
Meta CodeLlama 7B [21] 0.52 0.85 0.92 0.84 0.69 0.37 0.25 0.57 0.25 0.16
Google Gemma 2B [24] 0.51 0.66 0.76 0.74 0.67 0.42 0.28 0.55 0.30 0.27
- Avg. Performance 0.74 0.87 0.91 0.89 0.87 0.70 0.59 0.78 0.57 0.38
Table 1 shows both aggregate accuracy on MathCAMPS, as well as accuracy across standards partitioned by grade, whereas Figure 3 compares the aggregate accuracies on MathCAMPS and GSM8K. Closed-weights models are shown above the line, with open-weights models below. GPT-4o ranks at the top in overall accuracy. Since we used GPT-4 to generate the problems, we must rule out familiarity bias [22] in this result. We thus generated a 10%-scale dataset with the same pipeline but using Claude-3 Opus. We found that GPT-4o still outperforms Claude-3 Opus on this dataset (see Appendix B), suggesting that its advantage on MathCAMPS was not due to a familiarity bias. We make the following observations:
When considering overall performance, the trends in GSM8k hold on the novel problems from MathCAMPS, which cover overlapping topics (Pearson correlation of 0.865, $p<10^{-5}$; we show this correlation in Figure 3). This correlation corroborates the progress that public benchmarks have witnessed, suggesting that data contamination does not play a major role in explaining observed improvements in recent LLMs. We note that prior work attempting to replicate the distribution of GSM8k, such as the independent effort to collect GSM1k [26], has observed a smaller correlation, including substantial drops in performance for some models. This is entirely compatible with our findings here, due to the difficulty of exactly replicating the distribution over skills in any given human-created benchmark. As the sharp differences in Table 2 indicate, an (unintended) shift in this distibution can drastically — and unevenly — affect accuracy, even if no data contamination occurs. These shifts are easily avoided in an automated pipeline as in MathCAMPS, allowing us to draw new problems from the exact same distribution in the future.
Model Top outlier skill Rank change
GPT-4o 8.EE.C.8 - Solve two-variable systems ($1^{st}$ $22^{th}$)
Claude-3 Opus 2.MD.B.5 - Add/sub within 100 ($2^{nd}$ $13^{th}$)
Gemini-1.5 Pro K.OA.A.4 - Adding to equal 10 ($3^{rd}$ $19^{th}$)
Gemini-1.5 Flash 4.OA.B.4 - Factor pairs within 100 ($4^{th}$ $20^{th}$)
Claude-3 Haiku 3.OA.A.4 - Determine unknowns in mul/div probs ($9^{th}$ $1^{st}$)
Llama 3 70B K.OA.A.4 - Adding to equal 10 ($7^{th}$ $17^{th}$)
DeepSeek 67B K.NBT.A.1 - Decompose into 10s ($10^{th}$ $1^{st}$)
Llemma 34B K.OA.A.4 - Adding to equal 10 ($13^{th}$ $1^{st}$)
Mistral 7B 1.OA.A.1 - Add/sub within 20 ($14^{th}$ $21^{th}$)
DeepSeek Coder 33B 6.EE.A.1 - Evaluate exponents ($15^{th}$ $3^{rd}$)
Llemma 7B 6.EE.A.1 - Evaluate exponents ($18^{th}$ $5^{th}$)
Gemma 2B 8.EE.C.8 - Solve two-variable systems ($22^{th}$ $11^{th}$)
4.2 Follow-up tasks
Table 3: Model performance on our mathematical dialogue task, where the model must answer follow-up questions besides the initial problem. The second column, Accuracy with follow-ups, shows overall success rate across standards that contain follow-up questions, considering a model successful only when it answers a problem and its follow-up questions correctly. The third and fourth columns show the hardest standard for each model when it comes to follow-up questions, showing a standard’s code and abbreviated description, the model’s accuracy ignoring follow-ups, and after follow-ups.
Model Acc. with follow-ups Largest accuracy drop w/ follow-ups
GPT-4o 0.82 5.NF.A.1 - Add/sub fractions 0.90 0.61)
Claude-3 Opus 0.76 7.NS.A.1-fraction - Add/sub with fractions 0.57 0.25)
Gemini-1.5 Pro 0.77 5.NF.A.1 - Add/sub fractions 0.60 0.35)
Gemini-1.5 Flash 0.76 7.NS.A.1-fraction - Add/sub with fractions 0.78 0.38)
GPT-3.5 Turbo 0.71 7.NS.A.1-fraction - Add/sub with fractions 0.73 0.22)
Claude-3 Sonnet 0.72 5.NF.A.1 - Add/sub fractions 0.41 0.07)
Claude-3 Haiku 0.70 3.OA.A.3 - Mul/div within 100 1.00 0.73)
Llama 3 70B 0.69 4.NF.A.2 - Compare two fractions 0.99 0.66)
Mixtral 8x22B 0.69 7.NS.A.1-fraction - Add/sub with fractions 0.69 0.18)
DeepSeek 67B 0.68 6.NS.B.3 - Add/sub/mult/div decimals 0.59 0.37)
Llama 3 8B 0.58 4.NF.A.2 - Compare two fractions 0.90 0.52)
Mixtral 8x7B 0.58 5.NF.B.4 - Mult fractions 0.61 0.31)
Llemma 34B 0.55 5.NF.B.4 - Mult fractions 0.69 0.33)
Mistral 7B 0.48 7.NS.A.1-decimal - Add/sub with decimals 0.91 0.50)
DeepSeek Coder 33B 0.60 3.OA.A.3 - Mul/div within 100 0.95 0.81)
CodeLlama 34B 0.60 5.NF.B.4 - Mult fractions 0.52 0.39)
phi-2 0.39 7.NS.A.2 - Mult/div with fractions 0.57 0.08)
Llemma 7B 0.43 5.NF.B.4 - Mult fractions 0.61 0.22)
Gemma 7B 0.33 7.NS.A.1-decimal - Add/sub with decimals 0.91 0.32)
CodeLlama 13B 0.43 4.NBT.B.4 - Add/sub multi-digit nums 0.81 0.49)
CodeLlama 7B 0.49 2.NBT.B.7 - Add/sub within 100 0.80 0.67)
Gemma 2B 0.24 3.NBT.A.2 - Add/sub within 1000 0.93 0.26)
4.3 Learning dynamics
Finally, we use Pythia [6] to showcase another analysis that MathCAMPS enables: understanding the learning dynamics of mathematical skills during LM training. We evaluate checkpoints of Pythia 12B on all standards, and track the performance change as the model was trained. Figure 2 shows Pythia’s performance evolving during training on all 7 CC standards where the last checkpoint achieves at least 30% accuracy. Early in training, after 28k steps, Pythia performs best in a Kindergarten standard, K.OA.A.5 — “Fluently add and subtract within 5.”. At 57k steps, its performance is best in both K.OA.A.5 (37% accuracy) and two first-grade standards, 1.OA.A.1 and 1.OA.A.2 — both standards involve simple word problems with addition and subtraction within 20. Pythia starts to become proficient at a sixth-grade standard around midway during training: 6.EE.A.1, which involves evaluating simple expressions using whole-number exponents (e.g, computing squares and cubes). These skills develop in tandem with its linguistic competence – at first, Pythia repeats questions verbatim often, but at 57k steps it already often produces responses. Overall, the high-resolution of MathCAMPS as a reasoning benchmark can support future work to deepen our understanding of how language models acquire capabilities during training, and how specific factors (such as data, or scale) contribute to their learning.
Figure 2: Performance of Pythia 12B checkpoints on MathCAMPS standards as it evolves during training. We show all 7 standards where the last checkpoint has at least 30% accuracy.
@@ 5 Conclusion
We introduce MathCAMPS, a fine-grained synthetic benchmark of mathematical reasoning in LLMs. MathCAMPS is directly grounded on the Common Core Standards, a widely used curriculum in human education. By tying our problems to a human curriculum, we enable a much wider range of analyses to understand mathematical reasoning capabilities and weaknesses of LLMs. We show analyses of performance by grade level and identify particularly challenging skills for a range of models, though we believe these are only a few examples of analyses that MathCAMPS permits.
We note that MathCAMPS might also find applications in educational tools for human students, due to its correspondence to the Common Core. Future work in that direction will require psychometric analyses, to ensure that problem difficulty (aside from the abilities involved) is grade appropriate.
@@ Acknowledgments
Appendix A Common Core Standards in MathCAMPS
MathCAMPS is available on Github at https://github.com/gpoesia/mathcamps. All of the Common Core standards we implement are described in a configuration file, commoncore.yaml, where standards are instantiated by composing high-level components from the Common Core attribute grammar. Moreover, we provide our prompts used to generate the dataset and model responses, as well as all problems and model responses for all LLMs we evaluated.
We list the Common Core standards we represent in MathCAMPS in Tables 4 through 12, segregated by grade. Standards 3.MD.D.8, 4.MD.A.2, 7.NS.A.1, and 7.NS.A.3 are split up into sub-standards. This was done for ease of implementation of the grammar.

### v2

@@ 3.3 Generating follow-up questions
As human instructors know, follow-up questions are often a useful way to probe a student’s understanding. In MathCAMPS, we leverage our symbolic representation of problems to derive follow-up questions. We propose two kinds of questions: counterfactual questions, where we change a constant in the original problem, and incremental questions, where we add a new piece of information. For each CC standard, we mark which (if any) of these two categories of follow-ups are applicable. Symbolically, follow-up questions are represented as a difference to be applied to the original question — when we apply the difference, we obtain a new problem. We then use the same solver as the original problem to obtain the ground-truth answer to the follow-up question. We employ the same few-shot structure to translate this difference into a natural language question, and parse it back into a symbolic structure to test for cycle consistency.
4 Results
We use MathCAMPS to evaluate how the ability to reason mathematically evolves during pre- and post-training. We use open models that have released intermediate checkpoints (namely Amber from LLM360, OLMo2 7B from OLMo, and Pythia 12B from Eleuther AI). Then, we evaluate a subset of the available checkpoints on MathCAMPS. We use the evolution of checkpoint accuracy on the whole dataset and on subsets of the dataset to measure how different mathematical reasoning abilities evolve during training. Additionally, OLMo2 and Amber have instruction-tuned versions: we use these and Llama-3.1-8B’s versions, compared to their based pre-trained checkpoints, to study the impact of instruction tuning on individual Common Core standards. We also perform a comprehensive fine-grained (overall, per-grade, per-standard) evaluation of 23 popular open (including Llama 3, DeepSeek, Mistral of various sizes) and closed (e.g. GPT-4o, Claude 3, Gemini 1.5) models on MathCAMPS for which checkpoints are not available, and report all the results in the Appendix H. In the remainder of this section, we present our research questions and observations.
4.1 Pre-training
RQ1: How aligned are the learning trajectories of mathematical reasoning in LLMs compared to human students/curricula?
We first analyze how performance on skills from groups of grades evolve over time. Given overlap in the content for each grade, we split the grades K-8 into K-1, 2-3, 4-6, and 7-8. Then, we analyze how performance on these grade groups improves over time. We see the performance of the OLMo, LLM360, and Pythia models in Figure 2.
Figure 2: Model accuracy on problems coming from different grade groups evaluated during training. Each data point corresponds to an LLM checkpoint evaluated on MathCAMPS problems testing skills from the indicated range of grades. Training Progress (X-axis) is measured by percentage of total pre-training tokens seen by the checkpoint. Accuracy is final-answer accuracy on solving the problems with few-shot CoT prompting.
Notably, these models’ training data is not ordered according to any human curriculum. Despite this, we see that skills from earlier grades show higher performance earlier on in training, and this trend continues. It is possible that LLMs learn skills that are easier for humans first because these skills implicitly build on each other in the human reasoning data that LLMs observe. We see some evidence for this: when we look at the evolution of skills by grade groups, we see that the evolution of these groups follow the human curriculum, since on average, skills from earlier grades are also learned faster. However, even though the aggregate trend holds true, we also find that specific skills are learned in ways unaligned with the human curriculum. For example, 6.EE.A.1, a skill that asks students to evaluate numerical expressions involving whole-number exponents is learned very quickly, simply because calculating small exponents for small bases is an easily memorizable task. In contrast, even though human students could also learn these skills by memorization early on, they tend to learn them later only because they’re exposed to them later. Another important note is that we only evaluate the ability to solve ”computational” problems using these skills, whereas human students also need to pick up more conceptual skills (like explaining exponentiation, making analogies, illustrating numerical scenarious, etc). These abilities are not captured by simply memorizing the results of operations for small numbers.
Next, we analyze the relationship between model performance and the magnitude of the numbers involved in the problem. In Figure 3, we show the accuracy for each training checkpoint for the different numbers of digits seen in the answers. Interestingly, we note how model behavior does not follow exactly what is seen in humans. For example, even though problems with 1-digit answers are easier on average than those with two-digit answers, Amber seems to consistently over perform in cases where there are 2-digit answers. This holds true for the final Pythia model as well, but not for most other training checkpoints we looked at. This highlights a discrepancy that could potentially reveal underlying mechanisms that prevent LLMs from reasoning robustly and consistently.
Figure 3: Performance on problems of varying number of digits in their final answer, across pre-training checkpoints.
RQ2: Do skills emerge at a certain point in training, or do they evolve smoothly?
We evaluated each checkpoint on a diverse set of 49 skills across grades K-8. We observed that all skills tend to evolve smoothly during training, as opposed to certain skills disctretely emerging at a specific training step. In Figure 4, we show how all skills across grades 2 and 7 evolve across the Pythia, Amber, and OLMo models. Given space constraints, we include results for all other grades in Appendix B.
The granularity in MathCAMPS lets us conduct analyses that aren’t possible by looking at aggregate accuracies from datasets like GSM8K/MATH over time. For instance, the skills we included in 2nd grade mostly involve adding and subtracting within 100 under different settings (multi-step word problems, problems including currencies and lengths, etc.). Pythia demonstrates a particular performance gap between standards 2.OA.A.1 and 2.NBT.B.6. The prior encompasses the skill to solve one- and two-step problems with unknowns in all positions and the latter asks that students be able to add up to four two-digit numbers. A higher performance on 2.NBT.B.6 indicates that models performed well in a setting with more additions, but the linguistic complexity of 2.OA.A.1 made the problems harder (despite the numerical calculations in 2.OA.A.1 problems being easier). Amber and OLMo also demonstrate this performance gap in a smaller magnitude.
For seventh grade, however, there is more overlap in which skills models struggle with and which ones they are consistently performant on. 7.NS.A.1 is the skill about adding and subtracting rational numbers, with the fraction and decimal suffixes representing the fractional and decimal subsets of that standard. 7.NS.A.2 is about the multiplication and division of fractions, and 7.NS.A.3 involves solving real-world math problems that include the four operations and rational numbers. Interestingly, we see that across all models, skills involving operations with decimals are learned earlier on and tend to evolve faster. This is likely because operations with decimals are easier to memorize, given their high surface similarity to operations with integers (e.g. 0.28 + 0.45 is very similar to 28 + 45).
However, the skills involving fractions are learned at slower speeds, if they are learned at all. This is likely because operations with fractions require more reasoning steps per operation (e.g. calculating $\frac{7}{3}+\frac{2}{5}$ first requires finding a common denominator before conducting the easier step of addition). An alternate explanation involves the tokenization of numbers: it is likely that decimals and fractions are tokenized differently, and that the tokenization of decimals is more similar to that of integers. This would enable an easier transfer of mathematical reasoning skills learned for integers to skills applied to decimals. Additionally, we hypothesize that in some models, skill 7.NS.A.2 is learned relatively quickly because multiplying and dividing fractions requires fewer reasoning steps than adding and subtracting would.
(a) Grade 2 Refer to caption
(b) Grade 7 Refer to caption
Figure 4: Learning dynamics of individual Common Core standards in grades 2 and 7. Full results for all grades can be found in Appendix B.
Figure 5: Performance for models during training when also asked to answer follow-up questions about each problem. Here, we only consider problems that have at least one associated follow-up question (counterfactual or incremental).
These types of fine-grained analyses enabled by MathCAMPS might inform model pre-training and evaluation, especially for models that are trained with specific end uses in mind (e.g., in educational applications).
Figure 6: Effect of instruction tuning on individual skills in OLMo and LLM360. We show the difference in accuracy between the base model and the post-trained model: a negative value means that, on that particular standard, post-training hurts performance, whereas a positive value indicates that it improved. We find widely different profiles between the two instruction tuning methods: while the majority of skills in LLM360 improve, many more are either unaffected or hurt by the instruction tuning method applied to OLMo.
RQ3: Do models learn to apply reasoning skills in a robust manner?
We evaluate whether models’ reasoning skills are robust (and not a result of surface-level reasoning) by asking follow-up questions about each problem, probing for deeper understanding (Section 3.3 explains how we generate such questions). Figure 5 shows performance on problems and follow-up questions as training progresses. We observe that generally, as model ability to solve problems improve, so does their ability to answer follow-up questions about the problems they solve, demonstrating a relatively robust capability increase. However, the gap between the ability to answer one-turn and follow-up questions tends to increase over time. It is likely that this isn’t caused by a worsening ability to reason deeply, but rather because models solve more and more challenging problems as they train, and problems that are challenging are likely to have follow-ups that are also harder. Overall, we find a surprisingly robust ability to reason about mathematical problems developing during LLM training: not only the problems we use are novel, the models are increasingly capable of answering subsequent questions about those problems — a task we do not generally see in existing mathematical reasoning datasets, nor in post-training datasets focused on conversations (e.g., UltraChat (Ding et al., 2023)).
4.2 Post-training
RQ4: How does instruction tuning impact reasoning abilities?
Here, we use the Amber, OLMo, and Llama models to show how instruction tuning impacts specific mathematical skills. In Figure 6, we highlight Amber and OLMo. We hypothesize that the impact of instruction tuning on a model’s overall reasoning ability depends on the base model’s reasoning ability. We see that OLMo’s reasoning abilities were impacted more negatively as compared to Amber. This pattern of predominantly negative impact also holds for Llama-3.1-8B (Figure 7 in the Appendix). We note two potential reasons for this observation. First, the base Amber model is weaker than base OLMo and Llama, which means that there is more potential for reasoning abilities in Amber to develop or surface in response to instruction tuning. Second, for OLMo specifically, pre-training happened in two stages with distinct data distributions, with the first stage including 3.9T tokens of broad web-scale data and the second stage including 843B tokens of focused, high-quality mathematical and scientific data. Given that the second pre-training stage included high quality reasoning data, it’s likely that instruction tuning reversed some of the gains stemming from this second stage of training.
These results are also in alignment with an observation made by the concurrent work of Springer et al. (2025), which measures the impact of the quantity of pre-training data on performance changes due to fine-tuning. They show that the more tokens an LLM sees during pre-training, the less effective instruction tuning is on it. In our case, OLMo and Llama, which were trained on 15T and 4T tokens during pre-training respectively, suffered more degradation compared to Amber, which was only trained on  1.3T tokens during pre-training.
Although there is a broad variety in how specific skills are impacted by instruction tuning, we note a few interesting trends. First, we see how skills that are extremely challenging for the base model tend to stay unimpacted by instruction tuning. For example, 8.EE.C.8, a skill that measures the ability to solve systems of two equations in two variables, sees no improvement across all three models we inspect. Additionally, we also see some correlation in which skills improved or were hurt by instruction tuning across models. For example, in Figure 8, we see that skills that show improvements after instruction tuning tend to rely on multi-step reasoning, which possibly emerges due to better language understanding and instruction following. Overall, we find that instruction tuning has a variable impact on mathematical reasoning abilities of LLMs, with a nuanced interaction between model, specific skill, and the impact derived from post-training.
@@ 5 Conclusion
We introduce MathCAMPS, a fine-grained synthetic benchmark of mathematical reasoning in LLMs. MathCAMPS is directly grounded in the Common Core Standards, a widely used curriculum in human education. By tying our problems to a human curriculum, we enable a much wider range of analyses to understand how mathematical reasoning abilities evolve. We show analyses of alignment between human and LLM learning, the evolution of skill learning over training, robustness of reasoning skills over training, and the impact of instruction tuning on specific skills, though we believe these are only a few examples of analyses that MathCAMPS permits. MathCAMPS can also be used to directly compare how humans and LLMs acquire skills over time by measuring the performance of human participants of various ages on MathCAMPS.
@@ Acknowledgments
Appendix A Instruction-Tuning Results
In Figure 7, we see the results for instruction tuning on Llama-3.1-8B and in Figure 8, we see a side-by-side comparison of the impact of instruction tuning on all skills.
Figure 7: The impact of instruction-tuning on Llama 3.1 8B.
Figure 8: Impact (base minus post-trained performance) of instruction tuning across Amber, Llama, and OLMo, sorted by skill rather than by the performance delta. Overall, the direction of impact by skill tends to be similar across models, even if magnitude differs.
Appendix B Learning Dynamics for Grades K-8
Figures 9 through 17 show the learning dynamics for each grade for Amber, LLM360, and OLMo.
Figure 9: Learning Dynamics Across Amber, Pythia, OLMo for Kindergarten
Figure 10: Learning Dynamics Across Amber, Pythia, OLMo for First Grade
Figure 11: Learning Dynamics Across Amber, Pythia, OLMo for Second Grade
Figure 12: Learning Dynamics Across Amber, Pythia, OLMo for Third Grade
Figure 13: Learning Dynamics Across Amber, Pythia, OLMo for Fourth Grade
Figure 14: Learning Dynamics Across Amber, Pythia, OLMo for Fifth Grade
Figure 15: Learning Dynamics Across Amber, Pythia, OLMo for Sixth Grade
Figure 16: Learning Dynamics Across Amber, Pythia, OLMo for Seventh Grade
Figure 17: Learning Dynamics Across Amber, Pythia, OLMo for Eight Grade
Appendix C Common Core Standards in MathCAMPS
We will be releasing the code used to generate MathCAMPS and our first dataset. All of the Common Core standards we implement will be described in a configuration file, commoncore.yaml, where standards are instantiated by composing high-level components from the Common Core attribute grammar. Moreover, we provide our prompts used to generate the dataset and model responses, as well as all problems and model responses for all LLMs we evaluated.
We list the Common Core standards we represent in MathCAMPS in Tables 1 through 9, segregated by grade. Standards 3.MD.D.8, 4.MD.A.2, 7.NS.A.1, and 7.NS.A.3 are split up into sub-standards. This was done for ease of implementation of the grammar.